OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice

OpenStudy (anonymous):

oh youre doing flvs? mme too

OpenStudy (anonymous):

yeah you think you could help me?

OpenStudy (anonymous):

for the first equation you would try to solve for x

OpenStudy (anonymous):

and then plug that into the second equation

OpenStudy (anonymous):

soooo... x-y=16 x=16+y

OpenStudy (anonymous):

then substitute that into the second equation for the x variable

OpenStudy (anonymous):

so it would be 2(y+16)+y=2

OpenStudy (anonymous):

3y+32=2

OpenStudy (anonymous):

2y+32+y=2 3y+32=2 3y=-30 y=-10

OpenStudy (anonymous):

and the method is substitution

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

youre welcome!